# Humor in AI: Massive Scale Crowd-Sourced Preferences and Benchmarks for Cartoon Captioning

### v1

@@ 1 Introduction
This paper introduces a dataset and benchmark designed to investigate alignment in Large Language Models (LLMs). The dataset comprises a collection of over a quarter of a billion human ratings, curated from the New Yorker’s funny cartoon caption contest. The task of writing funny captions comes with significant challenges due to subjectivity and variability in human judgments. This benchmark represents a distinctive challenge for AI alignment, mirroring the complexities found in other tasks where expert humans consistently outshine state-of-the-art (SOTA) AI systems. Our exploration of alignment in LLMs through this dataset sheds light on fundamental questions regarding evaluation and alignment with human values and preferences.
Specifically, the paper studies the multifaceted realm of humor expression in LLMs, probing whether these models can effectively recognize humor and generate genuinely amusing captions resonating with human audiences. While LLMs may not have been explicitly tailored for humor, their exposure to diverse comedic content during training suggests an inherent capacity for humor recognition and expression within their computational repertoire. We propose a robust and scalable benchmark for evaluating a model’s humor capabilities that leverages advanced systems like GPT4o to potentially exceed human performance. This research aims to foster the development of more reliable, capable, and aligned AI systems, ultimately advancing the field of AI alignment and its practical applications.
@@
Our main contributions and findings are summarized as follows.
Dataset: The paper presents a pioneering dataset, consisting of massively crowdsourced ratings of human-generated cartoon captions from The New Yorker’s famous weekly cartoon caption contest. This dataset provides a unique opportunity to delve into the intricacies of humor generation in LLMs and marks the first large-scale dataset with human judgments for evaluating creative tasks. With over a quarter of a billion human ratings, this dataset offers a rich and diverse collection of examples, enabling researchers to explore the nuances of humor expression and perception in AI systems like never before. The dataset has already catalyzed groundbreaking research; a portion of it was shared with other researchers who utilized it in a recent award-winning paper at ACL 2023 [20]. By curating such a vast dataset, the paper not only facilitates deeper insights into the capabilities of LLMs but also lays the groundwork for future studies in this emerging field.
Benchmark: In addition to the dataset, the paper introduces novel metrics specifically tailored for evaluating the quality of humor generated by LLMs, leveraging cutting-edge techniques such as GPT. These metrics provide researchers with a standardized framework for assessing humor in AI-generated content, enabling precise and objective evaluations compared to traditional subjective methods. By establishing this benchmark, the paper not only enhances our understanding of humor in AI systems but also paves the way for more rigorous and systematic evaluations in the future.
Generative Capabilities of SOTA Models: Using the newly developed benchmark, the paper evaluates the performance of SOTA models such as GPT4o and Claude in generating humorous content relative to human-generated examples. By comparing the outputs of these models against human-generated humor, the paper offers valuable insights into the current capabilities and limitations of LLMs in the domain of humor generation. This analysis not only highlights areas where AI systems excel but also identifies areas for improvement, guiding future research efforts in enhancing the humor generation capabilities of LLMs.
Alignment Strategy Evaluation: Furthermore, the paper utilizes the benchmark to assess the effectiveness of various alignment strategies, including Reinforcement Learning from Human Feedback (RLHF), Direct Preference Optimization (DPO), and Best-of-N sampling (BoN), in enhancing humor generation in LLMs. By systematically comparing the performance of these alignment strategies against baseline models, the paper sheds light on their relative strengths and weaknesses, informing the development of more effective approaches for aligning AI systems with human preferences and values. This analysis not only advances our understanding of alignment techniques but also provides practical insights for improving the humor generation capabilities of LLMs in real-world applications.
In summary, the paper’s findings and contributions significantly advance LLM capabilities in humor evaluation and generation by providing a comprehensive dataset, establishing a benchmark for evaluation, and offering insights into the performance of SOTA models and alignment strategies. By addressing these key aspects, the paper not only enhances our understanding of humor in AI systems but also lays the foundation for future research and development in this rapidly evolving field. We open-source our dataset and code as linked in Appendix A.
@@ 2 Related Work
Alignment of LLMs. Finetuning of LLMs has proved a critical step in aligning the behavior of pretrained models to downstream tasks. A standard pipeline is to first finetune the pretrained model via supervised fine-tuning (SFT)—to imitate expert demonstrations—followed by reinforcement learning from human feedback (RLHF) [13]—where a reward model is trained on human preferences, and then the SFT model is trained to maximize this reward via PPO [41]. This pipeline has been successfully applied for finetuning frontier models [63, 4, 35, 51], and has inspired a vast amount of follow-up work refining and extending the SFT [59, 60, 18, 32, 16] and RLHF [6, 14, 30, 45, 33, 48, 9, 11] methodologies. Direct preference optimization (DPO) methods [38] have recently emerged as a simpler yet still effective replacement to the RLHF paradigm. Instead of training a reward model and then optimizing this reward, DPO combines these steps by directly optimizing the SFT model on offline human preference data, and has inspired a variety of extensions [19, 3, 46, 40, 50, 57]. While the aforementioned works focus on finetuning on human feedback, a related line of works has sought to finetune on AI-generated feedback [56, 54, 27, 8, 12, 58, 5]. Despite extensive research into various fine-tuning methodologies, understanding their effectiveness remains nascent. While several studies have explored when and why different methods are most effective [17, 26, 53, 10, 61, 43], they primarily address standard tasks like reducing harmfulness and increasing helpfulness, and do not assess fine-tuning methods for tasks requiring creativity, the focus of this work.
@@ 3 New Yorker Caption Contest
Number of contest 365
@@ 5.1 Experimental Results
MLLMs vs LLMs. Surprisingly, language-only models such as the pretrained Mistral model outperform the multimodal LLaVa model that has access to the entire cartoon images. Similarly, for overall group comparison, GPT-4o is also preferred over GPT-4o with vision. One possible reason is that the training corpus for multimodal LLMs can be much less diverse than the training corpus for the LLM. For example, LLaVa is only trained on a small multi-modal instruction following dataset ($\sim$80K unique images) [29], whereas generic LLMs like Mistral or Llama are trained on much larger dataset.
@@
Example Generation. As shown in Table 5, we provide some generation samples. Indeed, we see that LLMs generally produce longer and more verbose captions than top human ones.
@@ B.2 Caption Evaluation
System: You are a judge for the new yorker cartoon caption contest. Your job is to find the funniest caption. User: In this task, you will see a cartoon first and two captions that were written about it then. The task is to choose which caption is funnier. I am going to show you five cartoons, corresponding captions and their answers first. In the end, for the last cartoon, answer with only one letter A or B, and nothing else. User: $<$Insert Cartoon Image$>$ User: For this example, the two captions are A: $<$Insert CaptionA$>$ B: $<$Insert CaptionB$>$. The answer is Assistant: $<$Insert Answer$>$ ……Repeat user/assistant for four more examples…… User: $<$Insert Cartoon Image$>$ User: Find the funniest caption for each group. Then choose the funnier group based on these funniest captions. Think step by step but finish the last line of your answer with only one letter A or B, and nothing else. A: $<$Insert Caption Group A$>$ or B: $<$Insert Caption Group B$>$
@@ Appendix E Additional Results
We benchmark the performance of different reward model as in Table 8. weqweasdas/RM-Mistral-7B and Eurus-RM-7B Instruct are the top two models with the highest reward ranking accuracy. We choose to use weqweasdas/RM-Mistral-7B because it generally achieves better ranking accuracy for various data settings that we experimented on.

### v2

@@ 1 Introduction
This paper presents a dataset and benchmark for investigating alignment in Large Language Models (LLMs). Our dataset contains over a quarter of a billion human ratings from the New Yorker’s cartoon caption contest. Writing funny captions presents significant challenges due to the subjectivity of humor and variability in human judgments. This benchmark offers a unique challenge for AI alignment, reflecting complexities found in tasks where expert humans consistently outperform current AI systems. Our study examines fundamental questions about aligning them to generate funny captions similar to the winning captions that are most highly rated by the New Yorker readers.
We explore humor expression in LLMs, investigating whether these models can recognize humor and generate amusing captions that resonate with human audiences. While LLMs are not specifically designed for humor, their training on diverse content suggests a potential for humor recognition and expression. We propose a benchmark for evaluating a model’s humor capabilities using advanced systems like GPT-4.
Our empirical analysis shows that current LLMs can generate humorous captions but significantly underperform compared to high-ranking human submissions in the New Yorker’s caption contests. Generating successful captions requires multiple advanced capabilities: understanding of cultural references, recognition of humor patterns, logical reasoning, systematic planning, and visual analysis. The multi-component nature of caption generation makes this benchmark an effective test of broad LLM capabilities. Progress in aligning LLMs for this task will require both advancing these individual capabilities and developing methods to integrate them effectively. This benchmark therefore provides a comprehensive integration test for LLM capabilities.
Our main contributions and findings are:
Dataset: We present a large-scale dataset of human-rated cartoon captions from The New Yorker’s weekly contest. Each week, the New Yorker hosts a contest with a new cartoon, where thousands submit their funny captions. Hundreds of thousands of ratings are collected for each contest, and the winning captions are determined by those receiving the highest ratings. This dataset enables researchers to explore humor generation in LLMs and represents the first large-scale dataset with human judgments for evaluating creative tasks. With over 250 million ratings, it offers diverse examples for studying humor expression and perception in AI systems.
Benchmark: We introduce new metrics for evaluating humor quality in LLM-generated content, using GPT4 and group based techniques. These metrics provide a standardized framework for assessing AI-generated humor. Our benchmark allows for systematic comparisons between human and AI-generated humor.
Evaluation of State-of-the-Art Models: We assess the performance of models such as GPT-4 and Claude in generating humorous content, comparing their outputs to human-generated examples. This analysis offers insights into the current capabilities and limitations of LLMs in humor generation, identifying areas of strength and potential improvement.
Alignment Strategy Analysis: We use our benchmark to evaluate various alignment strategies, including Reinforcement Learning from Human Feedback (RLHF), Direct Preference Optimization (DPO), and Best-of-N sampling (BoN). By comparing these strategies, we provide insights into their effectiveness in enhancing humor generation in LLMs and aligning AI systems with human preferences.
In summary, our paper advances LLM capabilities in humor evaluation and generation through a comprehensive dataset, a new evaluation benchmark, and analysis of model performance and alignment strategies. This work enhances our understanding of humor in AI systems and provides a foundation for future research in this field. We open-source our dataset and code as detailed in Appendix A.
@@ 2 Related Work
Alignment of LLMs. Finetuning of LLMs has proved a critical step in aligning the behavior of pretrained models to downstream tasks. A standard pipeline is to first finetune the pretrained model via supervised fine-tuning (SFT)—to imitate expert demonstrations—followed by reinforcement learning from human feedback (RLHF) [14]—where a reward model is trained on human preferences, and then the SFT model is trained to maximize this reward via PPO [44]. This pipeline has been successfully applied for finetuning frontier models [68, 4, 38, 54], and has inspired a vast amount of follow-up work refining and extending the SFT [63, 64, 20, 35, 18] and RLHF [6, 16, 33, 48, 36, 51, 9, 11] methodologies. Direct preference optimization (DPO) methods [41] have recently emerged as a simpler yet still effective replacement to the RLHF paradigm. Instead of training a reward model and then optimizing this reward, DPO combines these steps by directly optimizing the SFT model on offline human preference data, and has inspired a variety of extensions [21, 3, 49, 43, 53, 61]. While the aforementioned works focus on finetuning on human feedback, a related line of works has sought to finetune on AI-generated feedback [60, 57, 30, 8, 13, 62, 5]. Despite extensive research into various fine-tuning methodologies, understanding their effectiveness for creative tasks remains nascent. While several studies have explored when and why different methods are most effective [19, 29, 56, 10, 66, 46], they primarily address standard tasks like reducing harmfulness and increasing helpfulness, and do not assess fine-tuning methods for tasks requiring creativity, the focus of this work.
@@
Humor in LLMs. Several recent works have studied humor capabilities of large language models, in addition to the ones studying the New Yorker Caption Contest. Concurrent to our work, Zhong et al. [65] also studies humor in a multi-modal setting, focusing on a different humor game Oogiri. Their experiments also suggests that existing chain of thought techniques are insufficient for LLMs to generate and understand humor. Similarly, Jentzsch and Kersting [26] also shows GPT-3 still lacks humor abilities despite the good performance on other factual knowledge benchmarks. Furthermore, several works have focused on LLMs’ capabilities in understanding humor, including humor detection [15], puns [58], and humor explanation [12].
@@ 3 New Yorker Caption Contest
Number of contests 365
Number of cartoons 365
@@ 5.1 Experimental Results
MLLMs vs LLMs. Surprisingly, language-only models such as the pretrained Mistral model outperform the multimodal LLaVa model that has access to the entire cartoon images. Similarly, for overall group comparison, GPT-4o is also preferred over GPT-4o with vision. To further investigate this issue, we conducted more experiments and obtained the following results:
  1. 1.
    GPT4-Turbo as evaluator given GPT4o descriptions (as reported in Table 2). Accuracy: 67%.
  2. 2.
    GPT4-Turbo-vision as evaluator given cartoon and GPT4o descriptions. Accuracy: 60.5%.
  3. 3.
    GPT4-Turbo-vision as evaluator given a blank image and GPT4o descriptions. Accuracy: 61.5%.
For bullet points 2 and 3 above, we are running the exact same model, the only difference is that one has access to the cartoon + text description, while the other has access only to the text description, thus isolating the effect of the image on the generation quality. We find that the visual element integration into the LLMs is negatively biasing the model’s accuracy. This observation is also consistent for overall and best pick group comparisons. Since giving a blank image also hurts performance compared to bullet point 1, GPT4 without vision, it is unlikely that the performance of the vision model is dragged down by the visual understanding capabilities.
One possible reason for the above observation is that the training corpus for multimodal LLMs can be much less diverse than the training corpus for the LLM. For example, LLaVa is only trained on a small multi-modal instruction following dataset ($\sim$80K unique images) [32], whereas generic LLMs like Mistral or Llama are trained on much larger dataset. Overall, these findings suggest there is still much research to be done in better integrating multi-modal capabilities into large language models.
@@
Example Generation and Qualitative Analysis. As shown in Table 5, we provide some generation samples for the cartoon in Figure 2. Indeed, we see that LLMs generally produce longer and more verbose captions than top human ones. Moreover, we generate multiple additional captions with GPT-4o-vision and Claude-3-Opus for the cartoon in Figure 2. Below, we make a qualitative analysis around the shortcoming of these generated captions from SOTA LLMs.
  • Missing visual details resulting in LLMs generating out-of-context captions. As an example, GPT-4o-vision generated another caption of “So the alien abduction statistics were right. Malls are the prime hunting grounds!” This caption does not match the cartoon though, since the aliens look missing and worried. Their bodies look skinny and weak. In other words, they don’t seem to be here to hunt humans.
  • Many generated captions are forms of word/phrase modification and creation. An example of this caption is “Do they have a Black Hole Friday sale?”, also generated by GPT4o-vision. Another example from Claude-3-opus is “I don’t see ’Invasion Supplies’ listed anywhere…” Both of these captions are inventing new words and phrases to make the caption funny. While these can be somewhat funny, they are usually not rated highly by the New Yorker audiences.
  • Winning captions tend to appeal to readers with multiple interpretations through different lenses. LLMs currently lack the ability to produce such captions. The New Yorker’s editorial pick of the final caption was “Oh sure, now you look at a map.” While this caption makes fun of the aliens blaming each other, it also references the experience of a driver missing the direction but claiming they knew the way. On the contrary, GPT4o-vision generated the caption “We travel light-years, and we still need directions!”, which despite making fun of the characters in the cartoon with the same concept, lacks the relatability from an additional perspective.
Overall, we think the GPT4 models are capable of generating humorous content. However, to rank among the top requires much deeper understanding of cultural references and more steps of reasoning before arriving at a high quality caption.
@@ B.2 Caption Evaluation
System: You are a judge for the new yorker cartoon caption contest. Your job is to find the funniest caption. User: In this task, you will see a cartoon first and two captions that were written about it then. The task is to choose which caption is funnier. I am going to show you five cartoons, corresponding captions and their answers first. In the end, for the last cartoon, answer with only one letter A or B, and nothing else. User: $<$Insert Cartoon Image$>$ User: For this example, the two captions are A: $<$Insert CaptionA$>$ B: $<$Insert CaptionB$>$. The answer is Assistant: $<$Insert Answer$>$ ……Repeat user/assistant for four more examples…… User: $<$Insert Cartoon Image$>$ User: Find the funniest caption for each group. Then choose the funnier group based on these funniest captions. Think step by step but finish the last line of your answer with only one letter A or B, and nothing else. A: $<$Insert Caption A$>$ or B: $<$Insert Caption B$>$
@@ Appendix E Additional Results
We benchmark the performance of different reward model as in Table 8. It is worth noting that our goal here is to understand the effect of the ranking model for the downstream preference learning algorithms, thus we evaluate on the same dataset as in Equation 1 instead of the setting of Table 2. weqweasdas/RM-Mistral-7B and Eurus-RM-7B Instruct are the top two models with the highest reward ranking accuracy. We choose to use weqweasdas/RM-Mistral-7B because it generally achieves better ranking accuracy for various data settings that we experimented on.